OpenStudy (anonymous):

Can someone give me a few example of a few simple cubic equations? One using ax^3+bx^2+c+d=0 formula

OpenStudy (amistre64):

you can create yor own by multiplying out: (x+a)(x+b)(x+c)=0 for any integer values of a,b, and c

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Please explain

OpenStudy (anonymous):

Hello SmoothMath, can you help me? :/

OpenStudy (anonymous):

ax^3+bx^2+c+d=0 x is your variable. a, b, c, and d are just constants. Pick any number for each one of them, and you have a cubic, as long as you don't pick 0 for a. Here's an example. a=2 b=3 c=2 d=1 gives you 2x^3 + 3x^2 +2x + 1 =0 Another example a = 1 b=-12 c=0 d=25 gives x^3 -12x^2 +25 = 0

OpenStudy (anonymous):

oooo okay so is that the final answer or do you solve for 0?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

it is amazingly hard to solve for a cubic equation what amsitre was suggesting is to start with your equation in factored form, then multiply out, and you will know what the zeros are

OpenStudy (anonymous):

But I mean... it seems like all the problem is asking for is "name a cubic equation" Which doesn't even require that you solve it. Setting one up is ridiculously simple.

OpenStudy (anonymous):

Sorry I must have worded it wrong I meant solving a cubic equation

OpenStudy (anonymous):

what is an example of solving a cubic equation

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ah, yes. Well then you should do what Amistre told you to do. Pick 3 linear factors and then multiply those three factors together. The solutions to the linear factors will be the solutions to your new cubic equation.